OpenStudy (anonymous):

Find the measure of each interior angle of a regular pentagon. answers - 176°, 108°, 120°, 150°

OpenStudy (ranga):

The exterior angles of all polygons add up to 360 degrees. A regular pentagon has 5 equal exterior angles that add up to 360. So each exterior angle = 360 / 5 = 72 degrees. Each interior angle = 180 - 72 = 108 degrees.

OpenStudy (anonymous):

thank you so much :)

OpenStudy (ranga):

You are welcome.